OpenStudy (anonymous):

Verify the identity sin 4x= 2 sin 2x cos 2x

OpenStudy (anonymous):

Can u help me see what 3\11 is equivalent to

OpenStudy (anonymous):

What?

hartnn (hartnn):

do you know the relation, \(\sin (a+b)=...?\)

hartnn (hartnn):

in that you can plug in a=2x, b=2x

OpenStudy (anonymous):

sin (a+b)= sinAcosB + sinBcosA

OpenStudy (anonymous):

or would sin (2x +2x)= sin 4x

hartnn (hartnn):

sin (a+b)= sin a cosb + sinb cos a YES put a= 2x, b= 2x in this. sin (2x +2x)= sin 4x YES

hartnn (hartnn):

so left side become sin 4x what does right side become ?